OpenStudy (uselesscookies):

Verify that each equation is an identity. 1 + sin 2x = (sinx + cosx)^2

OpenStudy (butterflydreamer):

Start off with your left hand side. So we have : LHS = 1 + sin2x (Now use the fact that sin2x = 2 sinx cosx) = 1 + 2sinxcosx (Now use the fact that sin^2x + cos^2x = 1) = sin^2x + cos^2x + 2sinxcosx (Re-arrange the equation and we get this:) = sin^2x + 2sinxcosx + cos^2x From this step ... you need to recall this basic quadratic rule thing LOL a^2 + 2ab + b^2 = (a + b)^2 In your case, the a= sinx and b =cosx Soooo, = (sinx + cosx)^2 =RHS tadaa

OpenStudy (uselesscookies):

Thank you >o<

zepdrix (zepdrix):

I think expanding the right hand side woulda been wayyy easier :3 but maybe that's just me hehe

OpenStudy (uselesscookies):

My teacher told me specifically to do the left hand side for some reason. ^^'

zepdrix (zepdrix):

When verifying, you're only allowed to mess with `one side`. Doesn't usually matter which side you choose to fiddle with though. Teacher wanted left side though huh? >.< Oh boyyy

OpenStudy (butterflydreamer):

you're welcome @uselesscookies xD I think usually we use the left side but ON occasion we use the right side :)